# ReSem3D: Refinable 3D Spatial Constraints via Fine-Grained Semantic Grounding for Generalizable Robotic Manipulation

### v3

@@ III-D MLLM-driven Automated Modeling of TAMP
As depicted in Fig. 4, the Language Model Program (LMP) framework is built on a hierarchical recursive architecture [ref14], enabling structured invocation and multi-level prompt nesting. Based on the framework, we have designed and encapsulated seven LMP modules within ReSem3D, corresponding respectively to Task Planner, Constraint Extraction, Constraint Refinement, Subtask Execution, Preconditions Building, Cost Function Building, and PostConditions Building. Each module implements an independent prompt structure and interfaces with system components through standardized APIs, enabling the automated modeling for task and motion planning.
During the LMP Setup phase, dependency relationships among LMP layers are established and mappings to API interfaces are configured driven by the natural language instructions. Subsequently, Task Planner performs global scheduling and intent parsing, collaborating with Constraint Extraction and Subtask Execution to model spatial constraints and execute tasks. Finally, considering the reliance on semantic understanding in spatial constraint modeling, Constraint Refinement is introduced to implement two-stage modeling with recursive parameter updates fed back to Task Planner.
Within the Subtask Execution phase, each task is decomposed into a sequence of constrained action execution units. Specifically, initiation conditions, optimization objectives, and termination criteria for action execution are formulated respectively by Preconditions Building, Cost Function Building, and Postconditions Building. Subsequently, each action execution unit activates the MPPI Execution, establishing the communication loop between the client and the Isaac Gym server. The client transmits real-time joint states and perception-fused cost functions, while the server performs GPU-accelerated MPPI sampling and optimization to synthesize continuous joint-space velocity commands, which are transmitted back to the client, thereby enabling stepwise closed-loop control.
Figure 5: ReSem3D is a unified robotic manipulation framework for semantically diverse environments. It leverages the synergy between MLLMs and VFMs to construct semantics-driven, two-stage hierarchical 3D spatial constraints, which are mapped into real-time optimization objectives in joint space to enable closed-loop perception-action control.
In the course of task execution, preconditions and postconditions are monitored. Upon violation of the precondition, the system backtracks and restarts the previous subtask, whereas satisfaction of the postcondition triggers progression to the next subtask. Ultimately, ReSem3D establishes a closed-loop TAMP framework centered on semantic understanding, guided by hierarchical recursion, and grounded in constraint feedback.
@@ IV Experiment
In this section, we first detail the experimental setup. Subsequently, we evaluate the cross-platform generalization of ReSem3D in simulated environments. Then, we demonstrate the manipulation generalization in the semantically diverse real-world environments. Furthermore, we analyze the core mechanisms of ReSem3D, exhibiting the superiority to baseline methods. Finally, we identify the key factors contributing to execution failures.
@@ IV-A Experimental Setup
Hardware Configuration. As shown in Fig. 6, our real-world experimental platform is constructed around the 6-DOF Universal Robots UR5e collaborative robotic arm, equipped with a PGI-140-80 two-finger parallel gripper and an Intel RealSense D435i RGB-D camera. For simulation, we employ the Omnigibson [ref46] platform developed by Stanford University, and utilize both the UR5e and the 7-DOF Franka Emika Panda collaborative arms.
@@
Tasks and Metrics. As shown in Fig. 5, we designed 12 real-world tasks (6 chemical lab scenarios and 6 household scenarios) to evaluate generalization and stability under static conditions and dynamic disturbances. In simulation, we perform 10 tasks (6 chemical lab scenarios and 4 household scenarios) to evaluate generalization across different robotic platforms under static conditions. Furthermore, we designed three simulated and four real-world experiments to assess the effectiveness in visual grounding. All tasks were evaluated over 10 trials, and success rate was reported.
Baselines. We compare against three representative frameworks: VoxPoser [ref10], ReKep [ref9] and CoPa [ref2]. VoxPoser integrates large language and vision models to generate 3D value maps interpretable by robots, enabling closed-loop trajectory synthesis without predefined motion primitives. It leverages GPT-4 and a visual frontend for semantic grounding of task objects, supporting flexible manipulation across diverse tasks. ReKep fuses VLMs and MLLMs to infer semantic keypoint constraints for structured tasks and solves end-effector trajectories in real time via hierarchical optimization. CoPa leverages foundation vision-language models to propose a two-stage task execution process, performing task-oriented grasping via coarse-to-fine visual grounding in the first stage, and identifying task-relevant spatial and geometric constraints for robot planning in the second stage. All three frameworks deeply integrate large language and vision models, employing semantics-driven visual grounding and control optimization to achieve efficient and robust robotic manipulation.
@@
Household Scenario Success Rate
Task Robot Platform Voxposer Rekep ReSem3D (Ours)
Toggle switch UR5e 0/10 3/10 9/10
@@
Cut Cake with Knife Franka Panda 0/10 5/10 7/10
Total 20.0% 42.5% 70.0%
Chemical Lab Scenario
Pour reagent into dish UR5e 7/10 0/10 8/10
@@
Pick Up Pestle on Mortar UR5e 0/10 3/10 6/10
@@
Total 36.6% 10.0% 65.0%
TABLE II: Visual Grounding Success Rates in Simulation
Task Scale ReKep ReSem3D (Ours)
@@
70% 0/10 9/10
40% 0/10 9/10
Pick Up Pestle on Mortar 150% 8/10 8/10
70% 7/10 9/10
40% 1/10 9/10
VFM and MLLM. We utilize GPT-4o [4] from OpenAI as a callable MLLM and combine it with the visual foundation model FastSAM [ref5] to enable high-efficiency multimodal reasoning (prompt design is available at ReSem3D.github.io. During the dynamic disturbance phase, we incorporate the TAPNext model from Google DeepMind [ref47] together with real-time RGB frame input to achieve online point tracking at approximately 20 Hz. This facilitates real-time inference of evolving spatial constraints in dynamic environments, thereby enhancing the stability and robustness of task execution.
IV-B Evaluation of ReSem3D in Simulated Environments
Figure 7: Comparison between Rekep and ReSem3D (Ours) in Constraint Modeling. Each method has two columns presenting RGB and point cloud visualizations, and two rows for 150% and 40% object scales. The numerical labels in purple, green, and red indicate the initial 2D constraints extracted by MLLM on the toaster, vase, and pestle, respectively, while the colorful arrows indicate the refinement of geometric and positional 3D constraints.
@@ IV-B Evaluation of ReSem3D in Simulated Environments
As summarized in Table I, we systematically evaluate the generalization capability of ReSem3D in simulation across different robotic platforms and semantically diverse environments, comprising 4 household and 6 chemical lab tasks under static conditions. For the Toggle Switch task, the density peak in positional constraint refinement is replaced by the maximum height to enable more direct extraction of symmetric point pairs. The results demonstrate that ReSem3D exhibits strong task adaptability and cross-platform generalization under zero-shot conditions, achieving execution success rates of 70% in household and 65% in chemical lab scenarios, respectively—consistently outperforming the baseline methods. This performance is attributed to the proposed two-stage 3D spatial constraint model. Specifically, VoxPoser achieves competent performance in tasks where the 3D center point of the object mask spatially aligns with a feasible grasp point (e.g., napkin, reagent, and stirrer). However, the success rate drops markedly when such alignment is absent. Although Rekep consistently generates semantically meaningful keypoints across tasks, its effectiveness is limited by the mismatch between predicted and task-relevant keypoints. Notably, since depth information for transparent containers is available in Omnigibson, these objects are modeled as opaque (e.g., plastic bottles) in simulation. Despite differences from real-world visual characteristics, this scheme preserves consistency in constraint generation and manipulation strategies, ensuring rigor in evaluation.
TABLE III: Success Rates in Real-world Environments.
Task Scenarios Voxposer Rekep ReSem3D (Ours)
Household Scenario Stage Static Dist. Static Dist. Static Dist.
Take Napkin M1 3/10 2/10 5/10 4/10 7/10 5/10
Fold Towel M1, M2 0/10 0/10 2/10 1/10 5/10 3/10
Place Lid on Mug M1, M3 0/10 0/10 3/10 1/10 5/10 4/10
Toggle switch M1, M3 0/10 0/10 1/10 0/10 4/10 3/10
Insert Flower into Vase M1, M3 0/10 0/10 2/10 1/10 6/10 4/10
Throw Trash into Bin M1, M3 7/10 6/10 6/10 5/10 8/10 7/10
@@
Place Beaker on Stirrer M1 0/10 0/10 1/10 0/10 6/10 5/10
Pour Liquid with Beaker M1 0/10 0/10 1/10 0/10 7/10 6/10
Pick Test Tube Rack M1 1/10 0/10 1/10 0/10 6/10 5/10
Pick Up Pestle on Mortar M1, M2 0/10 0/10 1/10 0/10 5/10 4/10
Grip Stir Bar with Tweezers M1, M2 0/10 0/10 0/10 0/10 4/10 2/10
Weigh the bottle M1, M3 7/10 5/10 5/10 4/10 8/10 6/10
@@
TABLE IV: Visual Grounding Success Rates in Real World
Task CoPa ReSem3D (Ours)
@@
Grip Stir Bar with Tweezers 0/10 6/10
Total 2.5% 60%
For further analysis, we focus on the constraint modeling adopted by ReKep for open-structured containers and dual-ended semantic objects. As shown in Fig. 7, the semantic keypoints of such objects are typically extracted by the MLLM, either computing the centroid of multiple opening-edge points (e.g., the vase and toaster) or identifying functionally distinct endpoints (e.g., the narrower end of the pestle). However, constraint generation in ReKep is sensitive to keypoint spacing and considerably dependent on object scale: scale increases tend to produce redundant constraints that disrupt accurate recognition, while scale decreases often cause constraint merging, leading to the loss of semantic keypoints. By comparison, ReSem3D demonstrates significantly greater stability under scale variation, exhibiting enhanced robustness against scale perturbations. This advantage is substantiated in Table II, where ReSem3D consistently outperforms baseline methods in visual grounding success rates on the Vase, Toaster and Pestle scaled to 150%, 70%, and 40%. These results indicate that the method effectively alleviates the impact of scale variation in open-structured container and dual-ended semantic object manipulation, thereby enhancing the zero-shot generalization.
IV-C Evaluation of ReSem3D in real-world Environments
Figure 8: Comparison between Baselines and ReSem3D (Ours) in Constraint Modeling and Closed-Loop Execution. Tasks 1 to 4 correspond to the execution for the bin, tweezer, towel, and beaker. The pink marks indicate end-effector positions or grasped constraints, yellow marks denote extracted constraints, green marks represent refined constraints, red marks highlight non-executable constraints, and colorful arrows show constraint refinement.
@@ IV-C Evaluation of ReSem3D in real-world Environments
In this section, we systematically evaluate the generalization of ReSem3D in real-world environments through a set of semantically diverse manipulation tasks, including 6 household and 6 chemical lab scenarios. As shown in Table III, M1, M2, and M3 indicate part-level constraint extraction, geometric constraint refinement, and positional constraint refinement, respectively. Static denotes environments without external disturbances, whereas “Dist.” refers to dynamic disturbances from manually displacing objects during task execution. Compared with baselines, ReSem3D achieves significant task execution performance, with 58.3% (Static) and 43.3% (Dist.) in the household scenario, and 60% (Static) and 45% (Dist.) in the chemical lab scenario, demonstrating superior zero-shot generalization and enhanced robustness against dynamic disturbances. A comprehensive analysis is presented below.
@@
Specifically, household scenarios are characterized by rich semantic structures and modest constraint precision, which enable baselines to retain considerable generalization in constraint modeling and object localization. Voxposer is robust in tasks where the 3D center points of object masks (e.g., napkins, trash, bins) are directly graspable, consistent with its simulation performance. For other tasks, failure occurs when the localization results do not meet the grasping requirements. Additionally, Rekep effectively extracts rich semantic keypoints for MLLM interpretation. However, complex semantic structures and fine-grained constraints (e.g., locating towel corners, detecting raised switch edges) surpass its interpretative capabilities. In contrast, chemical lab scenarios are typified by sparse semantic structure and prevalence of transparent materials, which degrade depth perception and necessitate greater spatial precision. These conditions further constrain baseline generalization (e.g., localizing beaker center, identifying pestle’s narrow tip), causing a continued decline in success rates. Building on these insights, ReSem3D introduces a two-stage 3D spatial constraint model that integrates semantic understanding and spatial reasoning, achieving 1.7% (Static) and 3.3% (Dist.) improvement in task success rates, maintaining strong robustness in semantically diverse environments.
@@
To qualitatively analyze the advantages of ReSem3D, we focus on its performance in semantic constraint modeling and closed-loop execution efficiency, as shown in Fig. 8. It can be observed that the semantic constraints from VoxPoser mainly concentrate around the coarse center of objects, lacking fine-grained geometric awareness and operability. Additionally, the absence of point clouds for transparent objects in chemical scenarios (e.g., beaker in Task 4) presents a significant challenge to accurate 3D center localization. Furthermore, the 5Hz 3D value map update rate restricts the responsiveness of closed-loop trajectory planning. Similarly, while ReKep constructs semantically directed constraints, it is limited in the accuracy (e.g., the center of the beaker, bin, and tweezer) and adjustability (e.g., the corners of the towel, the tips of the tweezer) of constraint generation. Its performance depends on the effectiveness of constraint generation and high-level reasoning by MLLM. Moreover, closed-loop optimization at 10Hz is constrained by the efficiency of inverse kinematics solving, leading to unstable real-time responsiveness.
@@
Furthermore, we compare with CoPa, which also adopts a coarse-to-fine visual grounding framework. As shown in Table IV, we evaluate four representative tasks from Table III to examine whether the visual localization is sufficient for successful task execution. Experimental results show that ReSem3D achieves a 60% success rate, significantly outperforming CoPa. A detailed qualitative analysis in Fig. 9 shows that, while CoPa performs hierarchical grounding, its fine-grained localization precision is limited to the part level, making it insufficient for tasks requiring finer spatial reasoning.

### v4

@@ III-D MLLM-driven Automated Modeling of TAMP
As depicted in Fig. 4, the Language Model Program (LMP) framework is built on a hierarchical recursive architecture [liang2023cap], enabling structured invocation and multi-level prompt nesting. Within this framework, we encapsulated seven LMP modules, including Task Planner, Semantic Structuring, Affordance Refinement, Subtask Execution, Preconditions Building, Cost Function Building, and PostConditions Building. Each module implements an independent prompt structure and interfaces with system components through standardized APIs, enabling automated modeling of task and motion planning.
During the LMP Setup, dependency relationships among LMPs are established and mappings to APIs are configured driven by the natural language instructions. Subsequently, Task Planner performs global scheduling and intent parsing, collaborating with Semantic Structuring and Subtask Execution to model affordance targets and execute tasks. Finally, Affordance Refinement further refines these targets through geometric and positional flows with recursive feedback.
In the Subtask Execution phase, tasks are decomposed into constrained action execution units. Specifically, initiation conditions, optimization objectives, and termination criteria are formulated respectively by Preconditions Building, Cost Function Building, and Postconditions Building. Subsequently, each action unit activates the MPPI Execution, establishing the communication loop between the client and Isaac Gym server. The client transmits real-time joint states and perception-fused cost functions, while the server performs GPU-accelerated MPPI sampling and optimization to synthesize continuous joint-space velocity commands, which are transmitted back to the client, enabling stepwise closed-loop control.
Figure 5: ReSemAct is a unified robotic manipulation framework for semantically diverse environments. It leverages synergistic reasoning between MLLMs and VFMs, introducing Semantic Structuring and Affordance Refinement (SSAR) to enable fine-grained robotic manipulation.
In the course of task execution, preconditions and postconditions are monitored. Upon violation of the precondition, the system backtracks and restarts the previous subtask, whereas satisfaction of the postcondition triggers progression to the next subtask. Ultimately, ReSemAct establishes a closed-loop TAMP framework centered on semantic understanding, guided by hierarchical recursion, and grounded in constraint feedback.
@@ IV Experiment
In this section, we first detail the experimental setup. Subsequently, we evaluate the cross-platform generalization of ReSemAct in simulated environments. Then, we demonstrate the manipulation generalization in the semantically diverse real-world environments. Furthermore, we conduct ablation experiment to evaluate the effectiveness of SSAR strategy. Finally, we analyze the core mechanisms of ReSemAct and identify the key factors contributing to execution failures.
@@ IV-A Experimental Setup
Hardware Configuration. As shown in Fig. 6, our real-world experimental platform is built around the 6-DOF Universal Robots UR5e robotic arm, equipped with a PGI-140-80 two-finger parallel gripper and an Intel RealSense D435i RGB-D camera. For simulation, we employ the Omnigibson [li2023behavior1k] platform developed by Stanford University, and utilize both the UR5e and the 7-DOF Franka Panda collaborative arms.
Tasks and Metrics. As shown in Fig. 5, we designed 12 real-world tasks (6 chemical lab scenarios and 6 household scenarios) to evaluate generalization and stability under static conditions and dynamic disturbances. In simulation, we perform 10 tasks (6 chemical lab scenarios and 4 household scenarios) to evaluate generalization across different robotic platforms under static conditions. Furthermore, we designed three simulated and four real-world experiments to assess the effectiveness in semantic grounding. All tasks were evaluated over 10 trials, and success rate was reported.
Baselines. We compare against three representative frameworks: VoxPoser [huang2023voxposer], ReKep [huang2024rekep] and CoPa [huang2024copa]. VoxPoser integrates large language and vision models to generate 3D value maps interpretable by robots, enabling closed-loop trajectory synthesis without predefined motion primitives. It leverages GPT-4 and a visual frontend for semantic grounding of task objects, supporting flexible manipulation across diverse tasks. ReKep fuses VLMs and MLLMs to infer semantic keypoint constraints for structured tasks and solves end-effector trajectories in real time via hierarchical optimization. CoPa leverages foundation vision-language models to propose a two-stage task execution process, performing task-oriented grasping via coarse-to-fine visual grounding in the first stage, and identifying task-relevant spatial and geometric constraints for robot planning in the second stage. All three frameworks deeply integrate large language and vision models, employing semantic grounding and control optimization to achieve efficient and robust robotic manipulation.
@@
VFM and MLLM. We utilize GPT-4o [achiam2023gpt-4] from OpenAI as a callable MLLM and combine it with the visual foundation model FastSAM [zhao2023fastsam] to enable high-efficiency multimodal reasoning (prompt design is available at ReSemAct.github.io. During the dynamic disturbance phase, we incorporate the TAPNext model from Google DeepMind [zholus2025tapnext] together with real-time RGB frame input to achieve online point tracking at approximately 20 Hz. This facilitates real-time inference of evolving affordance targets in dynamic environments, thereby enhancing the stability and robustness of task execution.
@@
Household Scenario Robot Platform Voxposer Rekep ReSemAct (Ours)
Toggle switch UR5e 0/10 3/10 8/10
@@
Cut Cake with Knife Franka Panda 0/10 5/10 6/10
Total 20.0% 42.5% 65.0%
Chemical Lab Scenario Robot Platform Voxposer Rekep ReSemAct (Ours)
Pour reagent into dish UR5e 7/10 0/10 7/10
@@
Pick Up Pestle on Mortar UR5e 0/10 3/10 5/10
@@
Total 36.6% 10.0% 61.6%
TABLE II: Semantic Grounding Success Rates in Simulation
Task Scale ReKep ReSemAct (Ours)
@@
70% 0/10 8/10
40% 0/10 8/10
Pick Up Pestle on Mortar 150% 7/10 7/10
70% 6/10 7/10
40% 1/10 8/10
IV-B Evaluation of ReSemAct in Simulated Environments
@@
As summarized in Table I, we systematically evaluate the generalization of ReSemAct in simulation across different robotic platforms and semantically diverse environments, comprising 4 household and 6 chemical lab tasks under static conditions. For the Toggle Switch task, the density peak in positional refinement Flow is replaced by the maximum height to enable more direct extraction of symmetric point pairs. The results demonstrate that ReSemAct exhibits strong task adaptability and cross-platform generalization, achieving execution success rates of 65% in household and 61.6% in chemical lab scenarios, respectively—consistently outperforming the baselines. This performance is attributed to the proposed SSAR. Specifically, VoxPoser achieves competent performance in tasks where the 3D center point of the object mask spatially aligns with a feasible grasp point (e.g., napkin, reagent, and stirrer). However, the success rate drops markedly when such alignment is absent. Although Rekep consistently generates semantically meaningful affordance targets across tasks, its effectiveness is limited by the mismatch between predicted and task-relevant affordance targets. Notably, since depth information for transparent containers is available in Omnigibson, these objects are modeled as opaque (e.g., plastic bottles) in simulation. Despite differences from real-world visual characteristics, this scheme preserves consistency in target generation, ensuring rigor in evaluation.
Figure 7: Comparison between Rekep and ReSemAct (Ours) in semantic grounding. Each method has two columns presenting RGB and point cloud visualizations, and two rows for 150% and 40% object scales. The numerical labels in purple, green, and red indicate the initial 2D constraints extracted by MLLM on the toaster, vase, and pestle, respectively, while the colorful arrows indicate the geometric and positional refinement.
For further analysis, we focus on the affordance targets adopted by ReKep for open-structured containers and dual-ended semantic objects. As shown in Fig. 7, the affordance targets of such objects are typically extracted by the MLLM, either computing the centroid of multiple opening-edge points (e.g., the vase and toaster) or identifying functionally distinct endpoints (e.g., the narrower end of the pestle). However, ReKep is sensitive to affordance targets spacing and considerably dependent on object scale: scale increases tend to produce redundant targets that disrupt accurate recognition, while scale decreases often cause targets merging, leading to the loss of affordance targets. By comparison, ReSemAct demonstrates significantly greater stability under scale variation, exhibiting enhanced robustness against scale perturbations. This advantage is substantiated in Table II, where ReSemAct consistently outperforms baseline methods in semantic grounding success rates on the Vase, Toaster and Pestle scaled to 150%, 70%, and 40%. These results indicate that the method effectively alleviates the impact of scale variation in open-structured container and dual-ended semantic object manipulation, thereby enhancing the fine-grained robot manipulation.
TABLE III: Success Rates in Real-world Environments
Task Scenarios Voxposer Rekep ReSemAct (Ours)
Household Scenario Static Dist. Static Dist. Static Dist.
Take Napkin 3/10 2/10 5/10 4/10 7/10 5/10
Fold Towel 0/10 0/10 2/10 1/10 5/10 3/10
Place Lid on Mug 0/10 0/10 3/10 1/10 5/10 4/10
Toggle switch 0/10 0/10 1/10 0/10 4/10 3/10
Insert Flower into Vase 0/10 0/10 2/10 1/10 6/10 4/10
Throw Trash into Bin 7/10 6/10 6/10 5/10 8/10 7/10
@@
Place Beaker on Stirrer 0/10 0/10 1/10 0/10 6/10 5/10
Pour Liquid with Beaker 0/10 0/10 1/10 0/10 7/10 6/10
Pick Test Tube Rack 1/10 0/10 1/10 0/10 6/10 5/10
Pick Up Pestle on Mortar 0/10 0/10 1/10 0/10 5/10 4/10
Grip Stir Bar with Tweezers 0/10 0/10 0/10 0/10 4/10 2/10
Weigh the Bottle 7/10 5/10 5/10 4/10 8/10 6/10
@@
TABLE IV: Semantic Grounding Success Rates in Real World
Task CoPa ReSemAct (Ours)
@@
Grip Stir Bar with Tweezers 0/10 5/10
Total 2.5% 57.5%
IV-C Evaluation of ReSemAct in real-world Environments
@@
In this section, we systematically evaluate the generalization of ReSemAct in the real world through semantically diverse manipulation tasks, including 6 household and 6 chemical lab scenarios. As shown in Table III, “Static” denotes environments without external disturbances, whereas “Dist.” refers to dynamic disturbances from manually displacing objects. Compared with baselines, ReSemAct achieves significantly superior performance, with 58.3% (Static) and 43.3% (Dist.) in the household scenario, and 60% (Static) and 46.6% (Dist.) in the chemical lab scenario, demonstrating enhanced generalization and robustness against dynamic disturbances. A comprehensive analysis is presented below.
@@
Specifically, household scenarios are characterized by rich semantic structures and modest manipulation precision, which enable baselines to retain considerable generalization in semantic grounding. Voxposer is robust in tasks where the 3D center points of object masks (e.g., napkins, trash, bins) are directly graspable, consistent with its simulation performance. For other tasks, failure occurs when the localization results do not meet the grasping requirements. Additionally, Rekep effectively extracts rich semantic affordance targets for MLLM interpretation. However, complex semantic structures and fine-grained targets (e.g., locating towel corners, detecting raised switch edges) surpass its interpretative capabilities. In contrast, chemical lab scenarios are typified by sparse semantic structure and prevalence of transparent materials, which degrade depth perception and necessitate greater spatial precision. These conditions further constrain baseline generalization (e.g., localizing beaker center, identifying pestle’s narrow tip), causing a continued decline in success rates. Building on these insights, ReSemAct introduces a Semantic Stucturing and Affordance Refinement Strategy that integrates fine-grained semantic grounding, achieving 1.7% (Static) and 3.3% (Dist.) improvement in task success rates, maintaining strong robustness in semantically diverse environments.
Figure 8: Comparison between Baselines and ReSemAct (Ours) in semantic grounding and closed-Loop execution. Tasks 1 to 4 correspond to the execution for the bin, tweezer, towel, and beaker. The pink marks indicate end-effector positions or grasped targets, yellow marks denote extracted course affordance anchor, green marks represent refined affordance targets, red marks highlight non-executable targets, and colorful arrows show target refinement.
@@
To qualitatively analyze the advantages of ReSemAct, we focus on its performance in semantic grounding and closed-loop execution efficiency, as shown in Fig. 8. It can be observed that the affordance targets from VoxPoser mainly concentrate around the object center, lacking fine-grained perception. Additionally, the absence of point clouds for transparent objects in chemical scenarios (e.g., beaker in Task 4) presents a significant challenge to accurate localization. Furthermore, the 5Hz 3D value map update rate restricts the responsiveness of closed-loop trajectory planning. Similarly, while ReKep constructs semantically directed targets, it is limited in the accuracy (e.g., the center of the beaker, bin, and tweezer) and refinement (e.g., the corners of the towel, the tips of the tweezer) of affordance targets. Its performance depends on the effectiveness of affordance generation and high-level reasoning by MLLM. Moreover, closed-loop optimization at 10Hz is constrained by the efficiency of inverse kinematics solving, leading to unstable real-time responsiveness.
@@
Furthermore, we compare with CoPa, which also adopts fine-grained robot manipulation. As shown in Table IV, we evaluate four representative tasks from Table III to examine whether the semantic grounding is sufficient for successful task execution. Experimental results show that ReSemAct achieves a 57.5% success rate, significantly outperforming CoPa. A detailed qualitative analysis in Fig. 9 shows that, while CoPa performs hierarchical grounding, its fine-grained localization precision is limited to the part level, making it insufficient for tasks requiring finer spatial reasoning.
IV-D Ablation Experiment
In this section, we conduct an ablation Experiment to evaluate the effectiveness of Semantic Structuring and Affordance Refinement (SSAR), as shown in Fig. 10. We perform semantic grounding on four representative tasks involving the tower, tweezer, beaker, and pestle. Specifically, w/o SSAR denotes a variant that directly relies on GPT-4o to predict affordance targets under language and visual prompts. w/o Semantic Affordance Description (SAD) replaces the affordance region with that of all objects, disrupting the semantic structure while maintaining normal operation. w/o AR removes the affordance refinement flow for reasoning.
The experimental results show that directly relying on MLLMs to predict affordance targets makes the tasks nearly infeasible. Similarly, once the semantic structure in SAD is disrupted, the system is unable to perform further fine-grained refinement. In addition, removing the affordance refinement has a limited impact on the Beaker task, as its affordance target largely coincides with the coarse-grained anchor. ReSemAct achieves the highest success rate across all tasks, demonstrating that SSAR plays a crucial role in fine-grained semantic grounding and significantly improves task success rates.